OpenStudy (anonymous):

on a trip mark drives the 600 mi to richmond averaging a certin speed. the return trip is made at an average speed that is 10mph slower total time for the round trip is 22hrs. find marks average speed on each part of the trip

OpenStudy (jamesj):

Call the speed for the trip to Richmond s. Then the time for that trip is, T1 = 600/s. Now, the speed for the second leg is s + 10. Find an expression now for T2. Then set T1 + T2 = 22 hours.

OpenStudy (anonymous):

t = d/r Solve the following for r:\[\frac{600}{r}+\frac{600}{r-10}=22\]r = 60 mph